I'm a school bus driver and want to know how to solve this problem handed to me from one of the students. -4x + 5y = 37. I honestly can't remember how to do this. It's been 35 years since math class for me. Sincerely, Sandra.
Answer by Alan3354(69443)

I'm a school bus driver and want to know how to solve this problem handed to me from one of the students. -4x + 5y = 37.
----------
It's 1 equation in 2 variables, x & y.
You can pick any value for x and find y, or vice versa.
----
There are an infinite # of solutions, no unique solution.
------
eg,
x = 0 --> y = 37/5
x = 1 --> y = 41/5
etc
--------------
If you graph it, you get a straight line of infinite length, and every point on the line is a "solution" of the equation.
------
-4x + 5y = 37
You can solve for x in terms of y:
-4x = -5y + 37
x = (5/4)y - 37/4
=====================
or y in terms of x
5y = 4x + 37
y = (4/5)x + 37/5